# HarmonicEval: Multi-modal, Multi-task, Multi-criteria Automatic Evaluation Using a Vision Language Model

### v1

@@ 4 Experiments
4.1 Settings
Baselines.
We implemented eight baseline metrics for the experiments on MMHE. Four of these are traditional n-gram-based metrics: BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), and CIDEr Vedantam et al. (2015). The remaining three metrics leverage neural network-based approaches: BERTScore Zhang et al. (2020), BARTScore Yuan et al. (2021), and CLIPScore Hessel et al. (2021). Additionally, we employ FLEUR Lee et al. (2024b), a VLM-based evaluation metric that provides a single overall score and its textual explanation. Details of these baselines are provided in Appendix E.
Performance measures.
@@ 4.1 Settings
To assess the performance of evaluation metrics on MMHE, we use accuracy (%) for overall evaluations and Kendall’s tau correlation coefficient $\tau$ for criterion-wise evaluations.
Implementation details.
The GPT-4o model is used as a backbone of HarmonicEval. The temperature hyperparameter is set to zero so that the outputs are as deterministic as possible for reproducibility. More details are in Appendix C.
4.2 Main results
Overall performance.
@@ 4.2 Main results
Table 2 compares HarmonicEval with eight metrics in terms of overall performance on MMHE. HarmonicEval achieves the highest accuracy on REG (66.6), VQA (76.4), VDU (73.4), and IC (77.0). While FLEUR also achieves the highest score of 76.4 on VQA, it falls short on VDU. These results underscore the effectiveness of HarmonicEval across multiple tasks.
Criterion-wise performance.
Table 3 shows the correlations between predicted scores and human judgments for each evaluation criterion. HarmonicEval evaluates numerical explainability by reporting correlations between its predicted criterion-wise scores and human judgment scores, whereas other metrics report correlations between their predicted overall scores and criterion-wise human judgment scores. As shown, HarmonicEval achieves the highest correlation across most criteria, highlighting its numerical explainability. FLEUR emerges as the second-best metric, a trend also observed in overall comparisons.
Task-wise trends in baseline metrics.
@@
To gain deeper insights into the evaluation behavior of baseline metrics, we examine the task-wise correlations. Specifically, we explore which evaluation criteria baseline metrics tend to prioritize or deprioritize for each task, potentially hindering comprehensive evaluation. In Table 3, black underlines denote the most correlated criterion, while gray underlines indicate the least correlated criterion for each metric within a task.
@@
For REG, completeness shows the highest correlation across most metrics. This is understandable, as REG requires explicit expressions to identify a unique object by distinguishing it from surrounding objects. For VQA, most metrics are more strongly correlated with conciseness but less so with completeness. This indicates that existing metrics deprioritize completeness, potentially leading to inaccurate evaluations of insufficient answers. A similar trend is observed in VDU, where completeness is also deprioritized by existing metrics. For IC, fluency exhibits low correlations for most metrics, suggesting a tendency to assign high scores even to texts lacking natural flow. A qualitative example illustrating this tendency is provided in Section 4.3.
4.3 Analysis.
@@
Qualitative examples.
We present a qualitative comparison between HarmonicEval and FLEUR. To generate textual explanations for HarmonicEval, we prompt the VLM with “Why? Tell me the reason.”, following Lee et al. (2024b).
Figure 5 illustrates two examples from the IC task. In Figure 5 (a), HarmonicEval detects severe fluency issues and reflects these in the overall score, whereas FLEUR fails to identify such deficiencies. This highlights that FLEUR may overlook specific evaluation criteria, such as fluency, as discussed in Section 1. In Figure 5 (b), HarmonicEval identifies errors in the caption more accurately than FLEUR, assigning a low correctness score and providing a clear, criterion-specific explanation. This suggests that HarmonicEval is able to offer not only numerical explainability but also textual explanation.
@@
w/o score smoothing
$\bm{67.5}$ $70.5$ $70.4$ $72.4$ $70.2$
Table 4: Ablation study. In the "w/o criterion-wise scoring" setting, we prompt the VLM to predict the overall score directly.
@@
VLM REG VQA VDU IC Avg.
LLaVA-7B
$62.9$ $72.0$ $61.4$ $73.4$ $67.4$
LLaVA-13B
$64.8$ $\bm{77.9}$ $63.2$ $72.9$ $69.7$
GPT-4o
$\bm{66.6}$ $76.4$ $\bm{73.4}$ $\bm{77.0}$ $\bm{73.4}$
Table 6: Overall accuracy (%) for different VLMs.
Ablation study.
@@
Table 4 highlights the effectiveness of three key components in HarmonicEval: criterion-wise scoring, score smoothing, and harmonic weighting. As shown, omitting criterion-wise scoring reduces performance on REG, VQA, and IC, indicating that comprehensive evaluation improves overall accuracy. Similarly, omitting harmonic weighting or score smoothing lowers performance across most tasks, demonstrating that these techniques stabilize the evaluation process.
VLM REG VQA VDU IC
Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc
LLaVA-7B
$14.5$ $14.0$ $14.8$ $\bm{28.4}$ $2.9$ $42.2$ $49.4$ $23.7$ $45.2$ $18.1$ $29.7$ $40.3$ $28.6$ $44.9$ $7.2$ $28.8$ $31.5$ $7.8$ $34.8$ $10.5$
LLaVA-13B
$\bm{31.4}$ $28.2$ $22.4$ $26.1$ $15.0$ $49.4$ $\bm{57.8}$ $26.7$ $49.7$ $30.5$ $39.7$ $45.9$ $36.4$ $42.2$ $27.7$ $35.6$ $40.0$ $4.5$ $27.4$ $5.2$
GPT-4o $23.2$ $\bm{30.8}$ $\bm{24.0}$ $20.7$ $\bm{23.8}$ $\bm{53.5}$ $50.6$ $\bm{31.8}$ $\bm{51.9}$ $\bm{44.4}$ $\bm{60.0}$ $\bm{48.8}$ $\bm{47.9}$ $\bm{51.2}$ $\bm{45.8}$ $\bm{44.7}$ $\bm{50.3}$ $\bm{19.8}$ $\bm{36.4}$ $\bm{22.8}$
Table 7: Criterion-wise performance for different VLMs.
Hyperparameter study.
Table 5 shows the hyperparameter study for $\gamma$. As expected, harmonic weighting using $1.00.5\leq\gamma<$ 1.0 performs better than uniform weighting ($\gamma=1.0$). Mathematical discussion underlying assumptions of these weightings is provided in Appendix A.
Backbone VLMs.
Tables 6 and 7 compare the overall performance and criterion-wise performance of three backbone VLMs: LLaVA-1.5-7B, LLaVA-1.5-13B, and GPT-4o. As shown, LLaVA-1.5-13B consistently perform better than LLaVA-1.5-13B. GPT-4o outperforms both models but leaves room for improvement, underscoring the need for future research to enhance alignment with human evaluations.
Metric Flickr8k-EX Flickr8k-CF
Reference-based BLEU 30.8 16.9
ROUGE 32.3 19.9
METEOR 41.8 22.2
CIDEr 43.9 24.6
SPICE 44.9 24.4
BERT-S 39.2 22.8
BERT-S++ 46.7
TIGEr 49.3
ViLBERTS-F 50.1
FAIEr-4 52.6 35.4
RefCLIP-Score 53.0 36.4
Polos $\bm{56.4}$ 37.8
RefFLEUR 51.9 38.8
Reference-free UMIC 46.8
FAIEr-r 50.1 32.4
CLIP-S 51.5 34.4
InfoCLIP 32.6 23.5
InfoMetIC 54.2 36.3
InfoMetIC+ 55.5 36.6
FLEUR 53.0 38.6
HarmonicEval $53.1$ $\bm{39.2}$
Table 8: Comparison on existing human evaluation datasets for IC task.
Performance on other IC datasets.
@@
To verify the generalizability of HarmonicEval’s effectiveness across different datasets, we additionally evaluated it on two human evaluation datasets: Flickr8k-EX and Flickr8k-CF. These datasets, widely used for image captioning, allow comparisons with a broader range of baselines, including learning-based metrics specifically designed for IC  [^2] We use Kendall’s tau-c and tau-b for Flickr8k-EX and Flickr8k-CF, respectively, following prior studies Hessel et al. (2021); Lee et al. (2024b) The details of additional baselines Anderson et al. (2016); Jiang et al. (2019); Lee et al. (2020); Hu et al. (2023); Wada et al. (2024) are in Appendix E..
@@
As shown in Table 8, HarmonicEval achieves performance comparable to or even better than state-of-the-art metrics on the three datasets. While our main focus is on automatic multi-task, multi-criteria evaluation, these results highlight the robustness of HarmonicEval across multiple datasets.
@@ 5 Related Work
Traditional metrics such as BLEU Papineni et al. (2002) and ROUGE Lin (2004) were originally proposed for automatic evaluation of machine translation and text summarization models. These metrics rely on n-gram overlap with reference texts. In recent years, large language models (LLMs) have been increasingly employed as evaluation metrics for various natural language generation (NLG) tasks Kocmi and Federmann (2023); Chiang and Lee (2023); Zheng et al. (2023); Song et al. (2024). For example, G-Eval Liu et al. (2023b) introduced a form-filling paradigm for criterion-based evaluation for text summarization and dialogue generation. Our HarmonicEval further extends this line of research with its capability of handling various vision-language tasks and provides a robust framework for aggregating criterion-wise scores based on the second-order statistics of token probability distributions assigned by VLMs.
For vision-language tasks, CIDEr Vedantam et al. (2015) measures the consensus between candidate and reference captions by weighting n-grams using TF-IDF. Recent metrics leverage VLMs to offer more flexible evaluation paradigms Lee et al. (2024a); Yu et al. (2024); Zhuang et al. (2024); Maeda et al. (2024); Chen et al. (2024). FLEUR Lee et al. (2024b), a state-of-the-art reference-free metric, provides textual explanations that underlie its overall scores. However, existing metrics primarily focus on providing overall evaluation, lacking numerical explainability and often overlooking specific criteria. HarmonicEval addresses these issues by offering criterion-wise scores alongside an overall score, enabling more interpretable and comprehensive evaluations.
@@
Several human evaluation datasets have been developed for NLG tasks. For instance, the WMT shared tasks Semenov et al. (2023) provide benchmarks for machine translation, TMU-GFM Yoshimura et al. (2020) for grammatical error correction, and SummEval Fabbri et al. (2021) for text summarization. In the vision-language domain, existing datasets such as Flickr8k Hodosh et al. (2013), PASCAL-50S Vedantam et al. (2015), Composite Aditya et al. (2018), THUMB Kasai et al. (2022), and Polaris Wada et al. (2024) target image captioning. In contrast, our MMHE dataset provides a multi-task, multi-criteria human evaluation resource, including sentences generated by several state-of-the-art VLMs across diverse vision-language tasks. This dataset enables more nuanced evaluations that capture a variety of criteria and tasks, advancing the field beyond the focus on captioning alone.
@@ 6 Conclusion
We introduced HarmonicEval and MMHE, a novel metric and dataset for automatic multi-criteria evaluation for various vision-language tasks, respectively. HarmonicEval is a reference-free evaluation metric that utilizes score variance to aggregate individual scores in a statistically reliable way. MMHE is a dataset covering four vision-language tasks and five criteria. Experimental results demonstrated the effectiveness of HarmonicEval on both MMHE and existing human evaluation datasets. We hope our study contributes to advancing research in automatic evaluation metrics.
@@
We carefully selected five general criteria that are considered effective across various vision-language tasks as detailed in Appendix B. These criteria were useful for discussions spanning the four vision-language tasks in this study. To further expand research to include a greater number of criteria and tasks, analyzing the relationship between task- or domain-specific criteria and these general criteria would also be necessary in future work.
@@ A.5 Harmonic weighting
We believe Assumption 3 is reasonless, and this justifies why $1.00.5<\gamma<$ 1.0 is used in practice.
@@ Appendix B Evaluation Criteria
We selected five key criteria essential for evaluating the output text of vision-language tasks: correctness, completeness, clarity, fluency, and conciseness.
@@
To validate these five criteria, we manually examined 100 instances across various vision-language tasks and confirmed their sufficiency for evaluating output text.
@@ C.1 Full prompts
Here, the underlined portions indicate task-dependent variations in the content. For example, “caption"" can be “answer"" for the VQA task. In Table 8, only the correctness and comprehensiveness prompts are used because datasets contain only short captions and for computational efficiency. In Tables 6 and 7, the prompts are slightly modified for each VLM to ensure that only the score is returned as the response.

### v2

@@ 4 Experiments
4.1 Performance on MMHE
We evaluate the performance of HarmonicEval on MMHE and compare it with conventional metrics. We also examine how existing metrics prioritize or deprioritize specific criteria in each task.
Settings.
@@ 4.1 Performance on MMHE
To assess the performance of each metric, we use accuracy (%) for the overall evaluation and the Kendall’s tau correlation coefficient $\tau$ for criterion-wise evaluations.
We implement eight baselines, including four n-gram-based metrics Papineni et al. (2002); Lin (2004); Vedantam et al. (2015); Banerjee and Lavie (2005) and four neural network-based metrics Zhang et al. (2020); Yuan et al. (2021); Hessel et al. (2021); Lee et al. (2024b), as listed in two groups in Table 2. Among them, FLEUR Lee et al. (2024b) is a state-of-the-art VLM-based metric. Note that all baseline metrics don’t support criterion-wise scoring and produce only overall scores. The details of baselines are in Appendix E. HarmonicEval utilizes GPT-4o as its backbone. For a fair comparison with FLEUR, we also implement GPT-FLEUR, which substitutes GPT-4o for the original LLaVA-1.5-13B.
Main results.
@@
Table 2 compares HarmonicEval with conventional metrics in terms of overall performance on MMHE. HarmonicEval achieves the highest accuracy in REG (66.6), VQA (76.4), IC (77.0), and attains the top average score of 73.4 across tasks. While GPT-FLEUR obtains the highest score on VDU (76.5), it performs less effectively on REG. These results underscore the strong multi-task capability of HarmonicEval.
Correlation analysis.
@@
To investigate the behavior of each existing evaluation metric, we show the correlation between the predicted overall scores and the human judgment scores for each criterion in Table 3. This analysis reveals which evaluation criteria conventional metrics tend to prioritize or deprioritize for each task, potentially hindering a comprehensive evaluation. In the table, red and blue underlines denote the most and least correlated criterion, respectively, for each task.
@@
We observe task-wise trends. For REG, completeness shows the highest correlation across most metrics. This is reasonable, as REG requires explicit expressions to identify a unique object by distinguishing it from surrounding objects. For VQA, most metrics are more strongly correlated with conciseness but less so with completeness. This indicates that conventional metrics deprioritize completeness, potentially leading to inaccurate evaluations of insufficient answers. A similar trend is observed in VDU, where completeness is also deprioritized by conventional metrics. For IC, fluency exhibits low correlations for most metrics, suggesting a tendency to assign high scores even to texts lacking natural flow. Overall, these results underscore the necessity of a comprehensive evaluation metric in multi-task scenarios.
Table 3 also shows the correlations between the criterion-wise scores of HarmonicEval and the human judgment scores. As shown, HarmonicEval achieves the highest correlation across most criteria, highlighting its ability for criterion-wise scoring.
4.2 Analysis
@@
Can HarmonicEval improve explainability?
Providing feedback on evaluation results to users is important. To investigate whether HarmonicEval offers clear textual explanations, we prompt the VLM with “Why? Tell me the reason.” after calculating the overall scores, following Lee et al. (2024b).
Figure 5 shows a qualitative comparison of HarmonicEval and FLEUR in the IC task. Panel (a) shows that HarmonicEval detects severe fluency issues and reflects this in its overall score, whereas FLEUR overlooks such deficiencies. This implies that FLEUR may miss certain criteria (e.g., fluency), as discussed in Section 3.3. Panel (b) demonstrates that HarmonicEval identifies incorrect details in the caption more precisely than FLEUR, assigning a low correctness score along with a criterion-specific explanation. These examples suggest that HarmonicEval can complement overall scores with informative textual feedback.
@@
Table 4: Ablation study.
@@
VLM Method REG VQA VDU IC Avg.
L-7B FLEUR $\bm{69.4}$ $69.1$ $\bm{63.2}$ $72.6$ $\bm{68.6}$
Harmonic $62.9$ $\bm{72.0}$ $61.4$ $\bm{73.4}$ $67.4$
L-13B FLEUR $62.9$ $76.4$ $60.9$ $\bm{73.9}$ $68.5$
Harmonic $\bm{64.8}$ $\bm{77.9}$ $\bm{63.2}$ $72.9$ $\bm{69.7}$
G-4o FLEUR $60.1$ $75.0$ $\bm{76.5}$ $76.0$ $71.9$
Harmonic $\bm{66.6}$ $\bm{76.4}$ $73.4$ $\bm{77.0}$ $\bm{73.4}$
Table 6: Comparison of HarmonicEval and FLEUR across different VLMs.
Is each component in HarmonicEval essential?
We conduct an ablation study for the two key components of HarmonicEval: criterion-wise scoring and harmonic weighting. Specifically, we examine two alternative approaches: 1) prompting the VLM to directly predict the overall score based on the five criteria, without using criterion-wise scores (w/o criterion-wise scoring); and 2) computing the overall score as a simple average of the criterion-wise scores (w/o harmonic weighing).
@@
Table 4 shows that both components improve overall performance. Removing criterion-wise scoring lowers scores on REG, VQA, and IC, indicating that explicit scoring for each criterion is vital for comprehensive evaluation. Similarly, omitting harmonic weighting reduces performance in most tasks, validating the effectiveness of our approach.
Is statistical aggregation effective?
Table 5 shows a hyperparameter study for $\gamma$. As expected, harmonic weighting with $\gamma=0.75$ performs the best. This justifies the importance of the statistical aggregation process.
Is HamonicEval effective across various backbone VLMs?
Table 6 shows the overall performance comparison between HarmonicEval and FLEUR using LLaVA-1.5-7B, LLaVA-1.5-13B, and GPT-4o as backbone models. HarmonicEval consistently outperforms FLEUR on both LLaVA-1.5-13B and GPT-4o. On the other hand, FLEUR slightly outperforms HarmonicEval when using LLaVA-1.5-7B. We found that this is because LLaVA-1.5-7B tends to underrate texts in the conciseness criterion compared to human-assigned scores. Nonetheless, these results highlight HarmonicEval’s effectiveness across different VLMs, particularly with more capable ones.
Metric F-EX F-CF Com Pas
Reference-based BLEU 30.8 16.9 30.6 72.9
ROUGE 32.3 19.9 32.4 74.1
METEOR 41.8 22.2 38.9 78.0
CIDEr 43.9 24.6 37.7 76.8
SPICE 44.9 24.4 40.3 69.6
BERT-S 39.2 22.8 30.1 79.1
TIGEr 49.3 45.4 80.7
ViLBERTS-F 50.1 52.4 79.6
FAIEr-4 52.6 35.4 57.7 81.4
RefCLIP-Score 53.0 36.4 55.4 83.1
Polos $\bm{56.4}$ 37.8 57.6 $\bm{86.5}$
RefFLEUR 51.9 38.8 64.2 85.5
Reference-free UMIC 46.8 30.1 56.1 85.1
FAIEr-r 50.1 32.4 50.5
CLIP-S 51.5 34.4 53.8 80.7
InfoCLIP 32.6 23.5 15.3 64.1
InfoMetIC 54.2 36.3 59.2 85.3
InfoMetIC+ 55.5 36.6 59.3 $\bm{86.5}$
FLEUR 53.0 38.6 63.5 83.2
GPT-FLEUR 53.5 39.0 61.5 82.6
HarmonicEval 53.1 $\bm{39.2}$ $\bm{66.2}$ 82.4
Table 7: Performance comparison on the Flickr8k-Expert (F-EX), Flickr8k-CF (F-CF), Composite (Com) and Pascal-50S (Pas) datasets.
Metric 1-ref 4-ref
Polos 93.3 95.4
RefFLEUR 97.3 $\bm{98.4}$
FLEUR 96.8 96.8
GPT-FLEUR 97.0 97.0
HarmonicEval $\bm{97.8}$ 97.8
Table 8: Performance comparison on the FOIL dataset.
4.3 Performance on conventional IC datasets
@@
To assess the robustness of HarmonicEval in standard image captioning scenarios, we also evaluate it on five widely used IC datasets: Flickr8k-Expert, Flickr8k-CF, Composite, Pascal-50S, and FOIL. These datasets allow comparisons with a broader set of baselines, including specialized IC metrics222 Following prior works Hessel et al. (2021); Lee et al. (2024b), we use Kendall’s tau-c for Flickr8k-EX and Composite, tau-b for Flickr8k-CF, and accuracy for Pascal-50S. Additional baselines Anderson et al. (2016); Jiang et al. (2019); Lee et al. (2020); Hu et al. (2023); Wada et al. (2024) are described in Appendix E..
@@
As shown in Table 7 and 8, HarmonicEval achieves performance comparable to or even better than state-of-the-art metrics on all five datasets. Although our primary goal is a multi-task evaluation, these results highlight HarmonicEval’s robustness even in single-task settings.
@@ 5 Related Work
Traditional metrics such as BLEU Papineni et al. (2002) and ROUGE Lin (2004) were originally proposed for the automatic evaluation of machine translation and text summarization models. These metrics rely on n-gram overlap with reference texts. In recent years, large language models (LLMs) have been increasingly employed as evaluation metrics for various natural language generation (NLG) tasks Kocmi and Federmann (2023); Chiang and Lee (2023); Zheng et al. (2023); Song et al. (2024). For example, G-Eval Liu et al. (2023b) introduced a form-filling paradigm for criterion-based evaluation for text summarization and dialogue generation. Our HarmonicEval further extends this line of research with its capability of handling various multi-modal tasks and provides a robust framework for aggregating criterion-wise scores based on the second-order statistics of token probability distributions assigned by VLMs.
For the IC task, CIDEr Vedantam et al. (2015) measures the consensus between candidate and reference captions by weighting n-grams using TF-IDF. Recent metrics leverage VLMs to offer more flexible evaluation paradigms Lee et al. (2024a); Yu et al. (2024); Zhuang et al. (2024); Maeda et al. (2024); Chen et al. (2024). FLEUR Lee et al. (2024b), a state-of-the-art reference-free metric, provides textual explanations that underlie its overall scores. However, existing metrics primarily focus on providing the overall evaluation, and often overlook specific criteria. HarmonicEval addresses these issues by offering criterion-wise scores alongside the overall score, enabling comprehensive evaluation.
@@
Several human evaluation datasets have been proposed for NLG tasks. Example datasets include SummEval Fabbri et al. (2021) for text summarization and the WMT shared tasks Semenov et al. (2023) for machine translation. In the multi-modal domain, existing datasets such as Flickr8k Hodosh et al. (2013), PASCAL-50S Vedantam et al. (2015), Composite Aditya et al. (2018), THUMB Kasai et al. (2022), and Polaris Wada et al. (2024) target image captioning. In contrast, our MMHE dataset provides a multi-task, multi-criteria human evaluation resource, including sentences generated by several state-of-the-art VLMs across diverse tasks. This dataset enables more nuanced evaluations that capture a variety of criteria and tasks, advancing the field beyond the focus on captioning alone.
@@ 6 Conclusion
We introduced HarmonicEval, a novel reference-free evaluation metric for multiple multi-modal tasks. HarmonicEval predicts criterion-wise scores and aggregates them via a statistically principled method to produce an overall score. In addition, we constructed MMHE, the first multi-task multi-criteria human evaluation dataset, consisting of 18,000 expert judgments. Experimental results demonstrate that HarmonicEval aligns more closely with human judgments than existing metrics on both MMHE and other commonly used human evaluation datasets. Furthermore, our analysis with MMHE reveals that existing metrics tend to prioritize certain criteria while neglecting others.
@@
We carefully selected five general criteria that are considered effective across various multi-modal tasks as detailed in Appendix B. These criteria were useful for discussions spanning the four multi-modal tasks in this study. To further expand research to include a greater number of criteria and tasks, analyzing the relationship between task- or domain-specific criteria and these general criteria would also be necessary in future work.
@@ A.5 Harmonic weighting
We believe Assumption 3 is reasonable, and this justifies why $1.00.5<\gamma<$ 1.0 is used in practice.
@@ Appendix B Evaluation Criteria
We selected five key criteria essential for evaluating the output text of multi-modal tasks: correctness, completeness, clarity, fluency, and conciseness.
@@
To validate these five criteria, we manually examined 100 instances across various multi-modal tasks and confirmed their sufficiency for evaluating output text.
@@ C.1 Full prompts
Here, the underlined portions indicate task-dependent variations in the content. For example, “caption"" can be “answer"" for the VQA task. In Table 7, only the correctness and comprehensiveness prompts are used because datasets contain only short captions and for computational efficiency. In Table 6, the prompts are slightly modified for each VLM to ensure that only the score is returned as the response.